# On Statistical Inference for High-Dimensional Binary Time Series

### v1

@@ 5.2 Consistency and Gaussian approximation
Lemma 17 discusses the Gaussian approximation theorem for the sample mean of dependent random variables. Gaussian approximation, as presented in Chernozhukov et al. (2013); Zhang and Wu (2017); Chernozhuokov et al. (2022); Chang et al. (2024a) among others is a useful tool for analyzing high-dimensional dependent data.
@@
Finally, we establish the consistency of the proposed bootstrap algorithm (Algorithm 4.2). According to Theorem 1.2.1 of Politis et al. (1999), the validity of Algorithm 4.2 is ensured if the cumulative distribution function of the bootstrap estimator in the bootstrap world converges to that of the original estimator. That is, the following condition holds true:
@@ 6.1 Simulation Study
Hyper-parameter selection. Several hyper-parameters are required to be selected in our approach, including the Lasso-type tuning parameter $\lambda$, the threshold $b_{d},$ and the bandwidth $h_{n}$ of Algorithm 4.2. While cross-validation is a common approach for hyper-parameter selection, randomly partitioning the data into training and testing sets destroys the data dependence structures in our setup, which, according to Burman et al. (1994), resists the cross-validation algorithms from selecting suitable hyper-parameters. Therefore, we propose to use a train-test split instead of commonly used cross-validations to tune parameters $\lambda$ and $b_{d},$ which are demonstrated as follows: We fix an split integer $n_{\mathtt{train}}<n$, and use $X_{1},\ldots,X_{n_{\mathtt{train}}}$ to fit the parameter $\widehat{\mathcal{A}}$. Based on the $n_{\mathtt{test}}$ data $X_{n_{\mathtt{train}}+1},\ldots,X_{n}$, we calculate the testing $L_{1}$ or $L_{2}$ norm error
@@ 6.2.1 Portfolio Management
After examining the mean performance, we fit a gbVAR(1) model to study both the temporal and cross-sectional dependence between data. Following the work of Jentsch and Reichmann (2022), we fit a preliminary Yule–Walker estimation to assess the regularity of data. It yields the largest row-sum in the parameter matrix $\max_{k}\sum_{l=1}^{K}|\widehat{\bm{\alpha}}_{kl}|=0.6712<1$. In addition, the largest absolute eigenvalue of $\widehat{\mathcal{A}}^{\mathtt{OLS}}$ is 0.1666, indicating that the underlying data-generating process of the stock prices is stationary and satisfies the assumptions required by our analysis.
@@
Compared to the baseline OLS estimator, our method shrinks small coefficients to zero. Furthermore, for a large proportion of the selected coefficients, the corresponding confidence intervals exclude zero, indicating their statistical significance. As shown in Figure 5(b), the state of Apple depends negatively on previous states of NVIDIA, and the state of Tesla also depends on both the states of Microsoft and Amazon the day before. This phenomenon can be explained by the ripple effect in the supply chain introduced in Kravchenko et al. (2024): the COVID pandemic in 2020-2021 triggered a global semiconductor shortage and a surge in demand for Tesla forced other companies such as Apple and Amazon to temper their expansion plans after the initial pandemic peak, which, in turn, affected their stock prices.
@@ 6.2.2 Global trade data
Another example concerns the annual international trade among $d=7$ countries (China, the United States, India, Japan, Germany, France, and the United Kingdom) from the year of 1950 to 2014 ($n=65$), with data obtained from Barbieri et al. (2009); Keshk (2017). We define an edge between two countries as 1 if the imports of country $i$ from country $j$ either begins or increases by more than 10%, and 0 otherwise. After collecting data, we fit a gbVAR(1) model with our proposed post-selection method, and demonstrate the result in Figure 6.
Figure 6 illustrates that the imports of China from Germany, France from China, India from France, and the United States from Germany retain the largest positive autoregressive coefficients on their own lags (ranging from 0.2 to 0.6) with strong persistence. The only positive cross‐regional spillover between Europe and the United States is along the import of the United States from Germany, with an estimated coefficient of about 0.3541. In Asia, the main bilateral trade flows between neighboring countries exhibit mutual positive spillovers, consistent with geographic proximity. The gbVAR model in this case features a highly sparse, modular network and represents the tightest internal connections and relatively point-to-point intercontinental connections. This modular network structure can be explained by the economic policy uncertainty within and through countries introduced by Tam (2018).
@@ 7 Conclusion
This manuscript extends the generalized binary vector autoregressive (gbVAR) model of Jentsch and Reichmann (2022) to a high-dimensional sparse binary time series setting, and adapts the post-selection estimator of Zhang (2023) for fitting the gbVAR model. To facilitate statistical inference, we propose a second-order wild bootstrap algorithm. Theoretically, we establish the asymptotic consistency and distributional properties of the post-selection estimator and prove the validity of the proposed bootstrap algorithm. Simulation and real-life data studies demonstrate that our method achieves good model selection and estimation performance relative to existing approaches, and has wide applications in modeling economic and trade data.
@@ B.3 Proof of Lemma 17
By Lemma 2 in Zhang and Politis (2023), we have a decomposition
@@
where $t=(1+\log(2d))/\psi$ and $h_{\psi,\psi,x}(\cdot)$ is well-defined in Zhang and Politis (2023). By Lemma 24, we have
@@
Similar to Lemma 7.3 from Zhang and Wu (2017) and (B.30) in Zhang and Politis (2023),
@@ B.5 Proof of Theorem 20
Now, we figure out $\max_{i=1,\ldots,d}|\mathcal{F}_{S_{i}}(\hat{\bm{\Sigma}}^{(0)\top}\hat{\bm{\Sigma}}^{(0)})-\mathcal{F}_{S_{i}}({\bm{\Sigma}}^{(0)\top}{\bm{\Sigma}^{(0)}})|_{1}$. By Horn and Johnson (2012), we have
@@ B.6 Proof of Corollary 21
Proof Algorithm 4.2 implies that $\bm{\Delta}^{*}_{ij}$ has joint normal distribution conditional on the observations $\{X_{t}\}$. From Lemma B.1 in Zhang and Politis (2023), we have for sufficiently large $n$,
@@ C.3 Proof of Lemma 27
Proof Suppose $Z_{t},1\leq t\leq n$, are martingale difference vectors with respect to the $\sigma$-field $\mathcal{G}_{t}$. By Theorem 4.1 of Pinelis (1994), we have
@@ C.4 Proof of Lemma 28
Proof If $d=1,2$, (47) can be obtained by Theorem 3.2 in Burkholder (1973) without much difficulty. If $d\geq 3$, for any vector $v=\left(v_{1},\ldots,v_{d}\right)^{\top}$, we have $|v|_{\infty}\leq|v|_{\log d}\leq d^{1/\log d}|v|_{\infty}=e|v|_{\infty}$. Hence, (47) is an immediate result of Lemma 27 by letting $s=\log d$.

### v2

@@ 5.2 Consistency and Gaussian approximation
Lemma 17 discusses the Gaussian approximation theorem for the sample mean of dependent random variables. Gaussian approximation, as presented in chernozhukov2013gaussian; zhang2017gaussian; chernozhuokov2022improved; chang2024central among others is a useful tool for analyzing high-dimensional dependent data.
@@
Finally, we establish the consistency of the proposed bootstrap algorithm (Algorithm 4.2). According to Theorem 1.2.1 of MR1707286, the validity of Algorithm 4.2 is ensured if the cumulative distribution function of the bootstrap estimator in the bootstrap world converges to that of the original estimator. That is, the following condition holds true:
@@ 6.1 Simulation Study
Hyper-parameter selection. Several hyper-parameters are required to be selected in our approach, including the Lasso-type tuning parameter $\lambda$, the threshold $b_{d},$ and the bandwidth $h_{n}$ of Algorithm 4.2. While cross-validation is a common approach for hyper-parameter selection, randomly partitioning the data into training and testing sets destroys the data dependence structures in our setup, which, according to burman1994cross, resists the cross-validation algorithms from selecting suitable hyper-parameters. Therefore, we propose to use a train-test split instead of commonly used cross-validations to tune parameters $\lambda$ and $b_{d},$ which are demonstrated as follows: We fix an split integer $n_{\mathtt{train}}<n$, and use $X_{1},\ldots,X_{n_{\mathtt{train}}}$ to fit the parameter $\widehat{\mathcal{A}}$. Based on the $n_{\mathtt{test}}$ data $X_{n_{\mathtt{train}}+1},\ldots,X_{n}$, we calculate the testing $L_{1}$ or $L_{2}$ norm error
@@ 6.2.1 Portfolio Management
After examining the mean performance, we fit a gbVAR(1) model to study both the temporal and cross-sectional dependence between data. Following the work of jentsch2022generalized, we fit a preliminary Yule–Walker estimation to assess the regularity of data. It yields the largest row-sum in the parameter matrix $\max_{k}\sum_{l=1}^{K}|\widehat{\bm{\alpha}}_{kl}|=0.6712<1$. In addition, the largest absolute eigenvalue of $\widehat{\mathcal{A}}^{\mathtt{OLS}}$ is 0.1666, indicating that the underlying data-generating process of the stock prices is stationary and satisfies the assumptions required by our analysis.
@@
Compared to the baseline OLS estimator, our method shrinks small coefficients to zero. Furthermore, for a large proportion of the selected coefficients, the corresponding confidence intervals exclude zero, indicating their statistical significance. As shown in Figure 5(b), the state of Apple depends negatively on previous states of NVIDIA, and the state of Tesla also depends on both the states of Microsoft and Amazon the day before. This phenomenon can be explained by the ripple effect in the supply chain introduced in kravchenko2024responding: the COVID pandemic in 2020-2021 triggered a global semiconductor shortage and a surge in demand for Tesla forced other companies such as Apple and Amazon to temper their expansion plans after the initial pandemic peak, which, in turn, affected their stock prices.
@@ 6.2.2 Global trade data
Another example concerns the annual international trade among $d=7$ countries (China, the United States, India, Japan, Germany, France, and the United Kingdom) from the year of 1950 to 2014 ($n=65$), with data obtained from barbieri2009trading; keshk2017correlates. We define an edge between two countries as 1 if the imports of country $i$ from country $j$ either begins or increases by more than 10%, and 0 otherwise. After collecting data, we fit a gbVAR(1) model with our proposed post-selection method, and demonstrate the result in Figure 6.
Figure 6 illustrates that the imports of China from Germany, France from China, India from France, and the United States from Germany retain the largest positive autoregressive coefficients on their own lags (ranging from 0.2 to 0.6) with strong persistence. The only positive cross‐regional spillover between Europe and the United States is along the import of the United States from Germany, with an estimated coefficient of about 0.3541. In Asia, the main bilateral trade flows between neighboring countries exhibit mutual positive spillovers, consistent with geographic proximity. The gbVAR model in this case features a highly sparse, modular network and represents the tightest internal connections and relatively point-to-point intercontinental connections. This modular network structure can be explained by the economic policy uncertainty within and through countries introduced by tam2018global.
@@ 7 Conclusion
This manuscript extends the generalized binary vector autoregressive (gbVAR) model of jentsch2022generalized to a high-dimensional sparse binary time series setting, and adapts the post-selection estimator of zhang2023statistical for fitting the gbVAR model. To facilitate statistical inference, we propose a second-order wild bootstrap algorithm. Theoretically, we establish the asymptotic consistency and distributional properties of the post-selection estimator and prove the validity of the proposed bootstrap algorithm. Simulation and real-life data studies demonstrate that our method achieves good model selection and estimation performance relative to existing approaches, and has wide applications in modeling economic and trade data.
@@ B.3 Proof of Lemma 17
By Lemma 2 in zhang2023debiased, we have a decomposition
@@
where $t=(1+\log(2d))/\psi$ and $h_{\psi,\psi,x}(\cdot)$ is well-defined in zhang2023debiased. By Lemma 24, we have
@@
Similar to Lemma 7.3 from zhang2017gaussian and (B.30) in zhang2023debiased,
@@ B.5 Proof of Theorem 20
Now, we figure out $\max_{i=1,\ldots,d}|\mathcal{F}_{S_{i}}(\hat{\bm{\Sigma}}^{(0)\top}\hat{\bm{\Sigma}}^{(0)})-\mathcal{F}_{S_{i}}({\bm{\Sigma}}^{(0)\top}{\bm{\Sigma}^{(0)}})|_{1}$. By horn2012matrix, we have
@@ B.6 Proof of Corollary 21
Proof Algorithm 4.2 implies that $\bm{\Delta}^{*}_{ij}$ has joint normal distribution conditional on the observations $\{X_{t}\}$. From Lemma B.1 in zhang2023debiased, we have for sufficiently large $n$,
@@ C.3 Proof of Lemma 27
Proof Suppose $Z_{t},1\leq t\leq n$, are martingale difference vectors with respect to the $\sigma$-field $\mathcal{G}_{t}$. By Theorem 4.1 of pinelis1994optimum, we have
@@ C.4 Proof of Lemma 28
Proof If $d=1,2$, (47) can be obtained by Theorem 3.2 in burkholder1973distribution without much difficulty. If $d\geq 3$, for any vector $v=\left(v_{1},\ldots,v_{d}\right)^{\top}$, we have $|v|_{\infty}\leq|v|_{\log d}\leq d^{1/\log d}|v|_{\infty}=e|v|_{\infty}$. Hence, (47) is an immediate result of Lemma 27 by letting $s=\log d$.